# Computational Complexity of Physical Information Sufficiency

### v3

@@ A.7 Complete Claim Coverage Matrix
cor:finite-budget-no-exact-admissibility Full DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_budget, DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement
cor:finite-budget-threshold-impossibility Full DQ.Physics.PhysicalCore.eventual_no_exact_admissibility_under_canonical_requirement, DQ.Physics.PhysicalCore.finite_budget_blocks_canonical_exponential_certification, DQ.Physics.PhysicalCore.finite_budget_blocks_exact_certification_beyond_threshold
cor:gap-externalization Full DQ.HardnessDistribution.saturatingSiteCost_eventually_constant, DQ.HardnessDistribution.totalDOF_ge_intrinsic
@@
cor:generalized-eventual-dominance Full DQ.HardnessDistribution.generalized_right_dominates_wrong_of_bounded_vs_identity_lower
cor:hardness-exact-certainty-inflation Full DQ.ClaimClosure.exact_certainty_inflation_under_hardness_core
cor:hardness-raw-only-exact Full DQ.ClaimClosure.exact_raw_only_of_no_exact_admissible_core, DQ.IntegrityCompetence.competence_implies_integrity
cor:information-barrier-query Full DQ.ClaimClosure.information_barrier_opt_oracle_core, DQ.ClaimClosure.information_barrier_state_batch_core, DQ.ClaimClosure.information_barrier_value_entry_core
cor:integrity-universal Full DQ.ClaimClosure.integrity_universal_applicability_core
@@
cor:linear-positive-no-saturation Full DQ.HardnessDistribution.native_dominates_manual
cor:no-auto-minimize Full DQ.ClaimClosure.no_auto_minimize_of_p_neq_conp
cor:no-uncertified-exact-claim Full DQ.ClaimClosure.no_uncertified_exact_claim_core
@@
cor:outside-excuses-no-exact-report Full DQ.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
@@
cor:physical-counterexample-core-failure Full CT.no_physical_counterexample_of_core_theorem, CT.physical_counterexample_invalidates_core_rule, CT.physical_counterexample_yields_core_counterexample, DQ.Physics.PhysicalCore.substrate_transport_blocks_encoded_counterexamples
cor:physics-no-universal-exact-claim Full DQ.ClaimClosure.no_exact_claim_admissible_under_hardness_core
cor:practice-bounded Full agent_transfer_licensed_iff_snapshot
@@
cor:practice-structured Full anchor_sigma2p_complete_conditional
cor:practice-tree Full anchor_sigma2p_reduction_core
cor:practice-unstructured Full ExactPhysicalClaimWellTyped
cor:query-obstruction-bool Full DQ.ClaimClosure.query_obstruction_boolean_corollary, RegimeSimulation
cor:right-wrong-hardness Full DQ.HardnessDistribution.no_positive_slope_linear_represents_saturating
cor:rlff-abstain-no-certs Full DQ.IntegrityCompetence.exactCertaintyInflation_iff_no_exact_competence, DQ.Physics.PhysicalCore.rlff_abstain_strict_global_maximizer_of_no_certificates
cor:type-system-threshold Full DQ.HardnessDistribution.linear_lt_exponential_plus_constant_eventually
@@
prop:bounded-slice-meta-irrelevance Full DQ.ClaimClosure.meta_coordinate_irrelevant_of_invariance_on_declared_slice, DQ.ClaimClosure.meta_coordinate_not_relevant_on_declared_slice
prop:bridge-failure-horizon Full DQ.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient, DQ.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency
prop:bridge-failure-stochastic Full DQ.ClaimClosure.standard_assumption_ledger_unpack
prop:bridge-failure-transition Full DQ.ClaimClosure.tractable_tree_core
prop:bridge-transfer-scope Full DQ.ClaimClosure.one_step_bridge
prop:budgeted-crossover Full DQ.ClaimClosure.physical_crossover_core, DQ.PhysicalBudgetCrossover.CrossoverAt
@@
prop:checking-witnessing-duality Full DQ.checking_witnessing_duality, DQ.checking_witnessing_duality_budget
prop:crossover-above-cap Full DQ.ClaimClosure.physical_crossover_above_cap_core, DQ.PhysicalBudgetCrossover.explicit_infeasible_succinct_feasible_of_crossover
prop:crossover-not-certification Full DQ.ClaimClosure.physical_crossover_hardness_core
prop:crossover-policy Full DQ.ClaimClosure.physical_crossover_policy_core
prop:decision-unit-time Full DQ.Physics.DecisionTime.decision_event_iff_eq_tick, DQ.Physics.DecisionTime.decision_taking_place_is_unit_of_time, DQ.Physics.DecisionTime.tick_increments_time, DQ.Physics.DecisionTime.tick_is_decision_event
prop:declared-contract-selection-validity Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core, DQ.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core, DQ.ClaimClosure.tree_structure_detectable
@@
prop:empty-sufficient-constant Full DP.anchorSufficient
@@
prop:generalized-assumption-boundary Full DQ.HardnessDistribution.generalizedTotal_with_saturation_eventually_constant, DQ.HardnessDistribution.generalized_dominance_can_fail_without_right_boundedness
prop:hardness-conservation Full DQ.HardnessDistribution.totalDOF_eventually_constant_iff_zero_distributed
prop:hardness-efficiency-interpretation Full DQ.HardnessDistribution.generalized_right_eventually_dominates_wrong
prop:heisenberg-strong-nontrivial-opt Full DQ.Physics.HeisenbergStrong.HeisenbergStrongBinding, DQ.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial, DQ.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty
prop:heuristic-reusability Full DQ.ClaimClosure.bounded_actions_detectable, DQ.ClaimClosure.reusable_heuristic_of_detectable, DQ.ClaimClosure.separable_detectable, DQ.ClaimClosure.transition_coupled_bridge_can_fail_on_sufficiency
prop:insufficiency-counterexample Full DP.classMonotoneOn, DP.constant_opt_all_sufficient
@@
prop:integrity-resource-bound Full DQ.ClaimClosure.integrity_resource_bound_for_sufficiency, DQ.IntegrityCompetence.completion_fraction_defined_of_declared_bound, DQ.IntegrityCompetence.evidence_nonempty_iff_claim_admissible
@@
prop:law-instance-objective-bridge Full CT.law_gap_physical_claim_holds, CT.no_law_gap_counterexample
@@
prop:one-step-bridge Full DQ.ClaimClosure.one_step_bridge
prop:oracle-lattice-strict Full ExplicitHardnessAssumptions, L.eval
prop:oracle-lattice-transfer Full DQ.ClaimClosure.oracle_lattice_transfer_as_regime_simulation, bounded_actions_detectable
prop:outside-excuses-explicit-assumptions Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core
@@
prop:physical-claim-transport Full CT.physical_claim_lifts_from_core, CT.physical_claim_lifts_from_core_conditional, DQ.Physics.PhysicalCore.substrate_transport_sound_conditional
prop:physics-no-universal-exact Full DQ.ClaimClosure.declared_physics_no_universal_exact_certifier_core
@@
prop:query-finite-state-generalization Full StandardComplexityAssumptions, adq_ordering
prop:query-randomized-robustness Full OneStepSequentialObjective, TwoStepObjective, adq
prop:query-randomized-weighted Full bridgeTransferLicensed, bridge_boundary_represented_family
prop:query-regime-obstruction Full DQ.ClaimClosure.query_obstruction_finite_state_core, StandardComplexityAssumptions, adq_ordering
prop:query-state-batch-lb Full StochasticCriterionObjective, agentBridgeClass
prop:query-subproblem-transfer Full DQ.ClaimClosure.regime_simulation_transfers_hardness, DQ.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency, DQ.ClaimClosure.subproblem_hardness_lifts_to_full
prop:query-tightness-full-scan Full TransitionCoupledObjective
prop:query-value-entry-lb Full StructureDetectable, boundaryCharacterized_iff_exists_sufficient_subset
prop:query-weighted-transfer Full bounded_actions_reusable_heuristic, bridgeFailureWitness
prop:raw-certified-bit-split Full DQ.ClaimClosure.certified_total_bits_split_core, DQ.IntegrityCompetence.admissible_irrational_strictly_more_than_rational, DQ.IntegrityCompetence.admissible_matrix_counts, DQ.IntegrityCompetence.certaintyInflation_iff_not_admissible, DQ.IntegrityCompetence.certificationOverheadBits, DQ.IntegrityCompetence.certificationOverheadBits_of_evidence, DQ.IntegrityCompetence.certificationOverheadBits_of_no_evidence, DQ.Physics.PhysicalCore.certified_work_lower_bound_raw
@@
prop:selector-separation Full DQ.ClaimClosure.selectorSufficient_not_implies_setSufficient
@@
prop:snapshot-process-typing Full DQ.ClaimClosure.agent_transfer_licensed_iff_snapshot, DQ.ClaimClosure.process_bridge_failure_witness, DQ.ClaimClosure.snapshot_vs_process_typed_boundary
@@
prop:sufficiency-char Full DQ.ClaimClosure.sufficiency_conp_reduction_core, DQ.ClaimClosure.sufficiency_iff_dq_ratio
prop:thermo-conservation-additive Full DQ.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
prop:thermo-hardness-bundle Full DQ.ClaimClosure.thermo_eventual_lift_core
prop:thermo-lift Full DQ.ClaimClosure.thermo_conservation_additive_core, DQ.ClaimClosure.thermo_energy_carbon_lift_core
prop:thermo-mandatory-cost Full DQ.ClaimClosure.thermo_hardness_bundle_core
@@
prop:typed-claim-admissibility Full DQ.ClaimClosure.tree_structure_detectable
@@
prop:universal-solver-framing Full DQ.ClaimClosure.typed_static_class_completeness
@@
thm:amortization Full DQ.HardnessDistribution.centralized_higher_leverage
thm:bridge-boundary-represented Full DQ.ClaimClosure.bridge_boundary_represented_family, DQ.ClaimClosure.bridge_failure_witness_non_one_step, DQ.ClaimClosure.bridge_transfer_iff_one_step_class
@@
thm:claim-integrity-meta Full DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core

### v4

@@ A.7 Complete Claim Coverage Matrix
cor:channel-degradation Full DQ.ClaimClosure.CH2, DQ.ClaimClosure.CH6
cor:ego-trap Unmapped (no derived Lean handle found)
@@
cor:finite-budget-no-exact-admissibility Full DQ.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split, DQ.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
cor:finite-budget-threshold-impossibility Full DQ.PhysicalBudgetCrossover.crossover_eventually_of_eventual_split, DQ.PhysicalBudgetCrossover.explicit_eventual_infeasibility_of_monotone_and_witness
cor:finite-lifetime Full DQ.ClaimClosure.SE5
cor:finite-state Full DQ.Physics.BoundedAcquisition.acquisition_rate_bound, DQ.Physics.BoundedAcquisition.one_bit_per_transition
cor:forced-finite-speed Full DQ.Physics.BoundedAcquisition.counting_gap_theorem
cor:gap-externalization Full DQ.HardnessDistribution.simplicityTax_grows, DQ.HardnessDistribution.totalExternalWork_eq_n_mul_gapCard
@@
cor:generalized-eventual-dominance Full DQ.HardnessDistribution.generalized_right_eventually_dominates_wrong
cor:ground-state Full DQ.Physics.BoundedAcquisition.srank_one_energy_minimum
cor:ground-state-passive Full DQ.ClaimClosure.AtomicCircuitExports.AC6
cor:hardness-exact-certainty-inflation Full DQ.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core
cor:hardness-raw-only-exact Full DQ.ClaimClosure.exact_certainty_inflation_under_hardness_core, DQ.IntegrityCompetence.competence_implies_integrity
cor:import-asymmetry Full DQ.ClaimClosure.SR4, DQ.ClaimClosure.SR5
cor:information-barrier-query Full DQ.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient, DQ.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency, DQ.ClaimClosure.information_barrier_opt_oracle_core
cor:integrity-universal Full DQ.ClaimClosure.information_barrier_value_entry_core
@@
cor:linear-positive-no-saturation Full DQ.HardnessDistribution.no_positive_slope_linear_represents_saturating
cor:neukart-vinokur Full DQ.ClaimClosure.process_bridge_failure_witness, DQ.ClaimClosure.selectorSufficient_not_implies_setSufficient
cor:no-auto-minimize Full DQ.ClaimClosure.minsuff_collapse_to_conp_conditional
cor:no-free-computation Full DQ.ClaimClosure.anchor_query_relation_true_iff, DQ.ClaimClosure.anchor_sigma2p_reduction_core, DQ.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset
cor:no-uncertified-exact-claim Full DQ.ClaimClosure.no_exact_claim_under_declared_assumptions_unless_excused_core
@@
cor:outside-excuses-no-exact-report Full DQ.ClaimClosure.no_auto_minimize_of_p_neq_conp
@@
cor:phase-transition Full DQ.ClaimClosure.IE12, DQ.ClaimClosure.IE13
cor:physical-counterexample-core-failure Full DQ.Physics.ClaimTransport.no_physical_counterexample_of_core_theorem, DQ.Physics.ClaimTransport.physical_counterexample_invalidates_core_rule, DQ.Physics.ClaimTransport.physical_counterexample_yields_core_counterexample
cor:physics-no-universal-exact-claim Full DQ.ClaimClosure.minsuff_conp_complete_conditional
cor:posed-anchor-certified-gate Full DQ.ClaimClosure.AQ8
cor:practice-bounded Full DQ.ClaimClosure.tractable_bounded_core
@@
cor:practice-structured Full DQ.ClaimClosure.tractable_separable_core
cor:practice-symmetry Full DQ.ClaimClosure.sufficiency_conp_complete_conditional
cor:practice-tensor Full DQ.ClaimClosure.subproblem_transfer_as_regime_simulation
cor:practice-tree Full DQ.ClaimClosure.tractable_tree_core
cor:practice-treewidth Full DQ.ClaimClosure.sufficiency_conp_reduction_core, DQ.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
cor:practice-unstructured Full DQ.ClaimClosure.hard_family_all_coords_core
cor:query-obstruction-bool Full DQ.ClaimClosure.physical_crossover_policy_core, DQ.ClaimClosure.process_bridge_failure_witness
cor:right-wrong-hardness Full DQ.HardnessDistribution.right_dominates_wrong
cor:rlff-abstain-no-certs Full DQ.IntegrityCompetence.exactCertaintyInflation_iff_no_exact_competence, DQ.IntegrityCompetence.rlff_abstain_strictly_prefers_no_certificates
cor:speed-integrity Full DQ.ClaimClosure.SE6
cor:theorem-equilibrium Full DQ.ClaimClosure.IE3, DQ.ClaimClosure.IE4, DQ.ClaimClosure.IE5
cor:thermo-dq Full DQ.ClaimClosure.DQ7, DQ.ClaimClosure.DQ8
cor:type-system-threshold Full DQ.HardnessDistribution.native_dominates_manual
cor:zero-capacity Full DQ.ClaimClosure.CH3
cor:zero-gap Full DQ.ClaimClosure.GE4, DQ.ClaimClosure.GE5
@@
prop:abstention-frontier Full DQ.ClaimClosure.no_auto_minimize_of_p_neq_conp
@@
prop:bounded-region Full DQ.Physics.BoundedAcquisition.BoundedRegion
prop:bounded-slice-meta-irrelevance Full DQ.ClaimClosure.integrity_resource_bound_for_sufficiency, DQ.ClaimClosure.integrity_universal_applicability_core
prop:bridge-failure-horizon Full DQ.ClaimClosure.explicit_state_upper_core, DQ.ClaimClosure.hard_family_all_coords_core
prop:bridge-failure-stochastic Full DQ.ClaimClosure.posed_anchor_query_truth_iff_exists_anchor
prop:bridge-failure-transition Full DQ.ClaimClosure.sufficiency_conp_reduction_core
prop:bridge-transfer-scope Full DQ.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
prop:budgeted-crossover Full DQ.ClaimClosure.oracle_lattice_transfer_as_regime_simulation, DQ.PhysicalBudgetCrossover.CrossoverAt
@@
prop:checking-witnessing-duality Full DQ.checkingBudgetPairs, DQ.checking_time_ge_witness_budget, DQ.witnessBudgetEmpty
prop:comp-thermo-chain Full DQ.ClaimClosure.DS5, DQ.ClaimClosure.DS6
prop:crossover-above-cap Full DQ.ClaimClosure.one_step_bridge, DQ.PhysicalBudgetCrossover.SuccinctInfeasible
prop:crossover-not-certification Full DQ.ClaimClosure.physical_crossover_above_cap_core
prop:crossover-policy Full DQ.ClaimClosure.physical_crossover_core
prop:decision-equivalence Full DQ.ClaimClosure.DE1, DQ.ClaimClosure.DE2, DQ.ClaimClosure.DE3, DQ.ClaimClosure.DE4
prop:decision-unit-time Full DQ.Physics.DecisionTime.decision_event_iff_eq_tick, DQ.Physics.DecisionTime.decision_event_implies_time_unit, DQ.Physics.DecisionTime.decision_taking_place_is_unit_of_time, DQ.Physics.DecisionTime.tick_is_decision_event
prop:declared-contract-selection-validity Full DQ.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core, DQ.ClaimClosure.no_auto_minimize_of_p_neq_conp, DQ.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
prop:discrete-state-time Full DQ.ClaimClosure.DS1, DQ.ClaimClosure.DS2
@@
prop:empty-sufficient-constant Full DQ.ClaimClosure.DP6
@@
prop:generalized-assumption-boundary Full DQ.HardnessDistribution.generalized_dominance_can_fail_without_right_boundedness, DQ.HardnessDistribution.generalized_dominance_can_fail_without_wrong_growth
prop:hardness-conservation Full DQ.HardnessDistribution.hardness_is_irreducible_required_work, DQ.HardnessDistribution.totalDOF_ge_intrinsic
prop:hardness-efficiency-interpretation Full DQ.HardnessDistribution.hardnessEfficiency_eq_central_share
prop:heisenberg-strong-nontrivial-opt Full DQ.Physics.HeisenbergStrong.strong_binding_implies_core_nontrivial, DQ.Physics.HeisenbergStrong.strong_binding_implies_nontrivial_opt_via_uncertainty, DQ.Physics.HeisenbergStrong.strong_binding_implies_physical_nontrivial_opt_assumption
prop:heuristic-reusability Full DQ.ClaimClosure.anchor_query_relation_true_iff, DQ.ClaimClosure.posed_anchor_checked_true_implies_truth, DQ.ClaimClosure.posed_anchor_exact_claim_requires_evidence, DQ.ClaimClosure.sufficiency_iff_dq_ratio
prop:identifiability-convergence Full DQ.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core
prop:insufficiency-counterexample Full DQ.ClaimClosure.DP7, DQ.ClaimClosure.DP8
@@
prop:integrity-prerequisite Full DQ.ClaimClosure.IE17
prop:integrity-resource-bound Full DQ.ClaimClosure.information_barrier_state_batch_core, DQ.IntegrityCompetence.completion_fraction_defined_of_declared_bound, DQ.IntegrityCompetence.evidence_nonempty_iff_claim_admissible
@@
prop:landauer-constraint Full DQ.ClaimClosure.IE1, DQ.ClaimClosure.IE6
prop:law-instance-objective-bridge Full DQ.Physics.ClaimTransport.PhysicalEncoding, DQ.Physics.ClaimTransport.physical_claim_lifts_from_core
@@
prop:lorentz-discrete Full DQ.ClaimClosure.DS3, DQ.ClaimClosure.DS4
prop:mdp-tractable Full DQ.ClaimClosure.process_bridge_failure_witness
@@
prop:one-step-bridge Full DQ.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
prop:oracle-lattice-strict Full DQ.ClaimClosure.horizonTwoWitness_immediate_empty_sufficient, DQ.ClaimClosure.information_barrier_opt_oracle_core
prop:oracle-lattice-transfer Full DQ.ClaimClosure.no_uncertified_exact_claim_core, DQ.ClaimClosure.pose_returns_anchor_query_object
prop:orbital-symmetry Full DQ.ClaimClosure.AtomicCircuitExports.AC8
prop:outside-excuses-explicit-assumptions Full DQ.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core
@@
prop:physical-claim-transport Full DQ.Physics.ClaimTransport.no_physical_counterexample_of_core_theorem, DQ.Physics.ClaimTransport.physical_claim_lifts_from_core_conditional, DQ.Physics.ClaimTransport.physical_counterexample_invalidates_core_rule
prop:physics-no-universal-exact Full DQ.ClaimClosure.declaredBudgetSlice
@@
prop:pose-anchor-object Full DQ.ClaimClosure.AQ1, DQ.ClaimClosure.AQ2, DQ.ClaimClosure.AQ3
prop:posed-anchor-typed-exact Full DQ.ClaimClosure.AQ4, DQ.ClaimClosure.AQ5, DQ.ClaimClosure.AQ6, DQ.ClaimClosure.AQ7
prop:query-finite-state-generalization Full DQ.ClaimClosure.process_bridge_failure_witness
prop:query-randomized-robustness Full DQ.ClaimClosure.process_bridge_failure_witness
prop:query-randomized-weighted Full DQ.ClaimClosure.process_bridge_failure_witness
prop:query-regime-obstruction Full DQ.ClaimClosure.process_bridge_failure_witness
prop:query-state-batch-lb Full DQ.ClaimClosure.horizon_gt_one_bridge_can_fail_on_sufficiency, DQ.ClaimClosure.process_bridge_failure_witness
prop:query-subproblem-transfer Full DQ.ClaimClosure.pose_returns_anchor_query_object, DQ.ClaimClosure.posed_anchor_query_truth_iff_exists_forall, DQ.ClaimClosure.posed_anchor_signal_positive_certified_implies_admissible
prop:query-tightness-full-scan Full DQ.ClaimClosure.process_bridge_failure_witness
prop:query-value-entry-lb Full DQ.ClaimClosure.information_barrier_opt_oracle_core, DQ.ClaimClosure.process_bridge_failure_witness
prop:query-weighted-transfer Full DQ.ClaimClosure.process_bridge_failure_witness
prop:raw-certified-bit-split Full DQ.ClaimClosure.bridge_failure_witness_non_one_step, DQ.IntegrityCompetence.admissible_irrational_strictly_more_than_rational, DQ.IntegrityCompetence.admissible_matrix_counts, DQ.IntegrityCompetence.certaintyInflation_iff_not_admissible, DQ.IntegrityCompetence.certificationOverheadBits, DQ.IntegrityCompetence.certificationOverheadBits_of_evidence, DQ.IntegrityCompetence.certificationOverheadBits_of_no_evidence, DQ.IntegrityCompetence.certifiedTotalBits_ge_raw, DQ.IntegrityCompetence.certifiedTotalBits_gt_raw_of_evidence
prop:reaction-competence Full DQ.ClaimClosure.MI3, DQ.ClaimClosure.MI4
prop:refinement-strengthens Full DQ.ClaimClosure.stochastic_objective_bridge_can_fail_on_sufficiency
prop:retraction-evidence-integrity Full DQ.ClaimClosure.subproblem_hardness_lifts_to_full
prop:retraction-no-evidence-violates Full DQ.ClaimClosure.subproblem_transfer_as_regime_simulation
@@
prop:selector-separation Full DQ.ClaimClosure.posed_anchor_exact_claim_admissible_iff_competent
@@
prop:sequential-anchor-refinement Full DQ.StochasticSequential.sequential_anchor_sufficient_of_sequential_sufficient
prop:sequential-anchor-tqbf-reduction Full DQ.StochasticSequential.SequentialAnchorCheckInstance, DQ.StochasticSequential.reduceTQBF_correct_anchor, DQ.StochasticSequential.reduceTQBF_to_sequential_anchor_reduction
prop:sequential-bounded-horizon Full DQ.ClaimClosure.physical_crossover_policy_core
prop:sequential-static-relation Full DQ.ClaimClosure.physical_crossover_above_cap_core
@@
prop:snapshot-process-typing Full DQ.ClaimClosure.physical_crossover_hardness_core, DQ.ClaimClosure.posed_anchor_no_competence_no_exact_claim, DQ.ClaimClosure.system_transfer_licensed_iff_snapshot
prop:static-stochastic-strict Full DQ.StochasticSequential.static_stochastic_strict_separation
prop:static-stochastic-transfer Full DQ.ClaimClosure.pose_returns_anchor_query_object
@@
prop:stochastic-anchor-refinement Full DQ.StochasticSequential.stochastic_anchor_sufficient_of_stochastic_sufficient
prop:stochastic-anchor-strict-reduction Full DQ.StochasticSequential.StochasticAnchorCheckInstance, DQ.StochasticSequential.reduceMAJSAT_correct_anchor_strict, DQ.StochasticSequential.reduceMAJSAT_to_stochastic_anchor_reduction
prop:stochastic-bounded-support Full DQ.ClaimClosure.posed_anchor_signal_positive_certified_implies_admissible
prop:stochastic-landauer-floor Full DQ.StochasticSequential.landauerEnergyFloor_mono_bits, DQ.StochasticSequential.landauerEnergyFloor_nonneg, DQ.StochasticSequential.thermodynamicCost_eq_landauerEnergyFloorRoom_states
prop:stochastic-potential-duality Full DQ.StochasticSequential.stochasticExpectedUtility_eq_neg_expectedActionPotential, DQ.StochasticSequential.stochasticExpectedUtility_le_iff_expectedActionPotential_ge, DQ.StochasticSequential.utilityFromPotentialDrop_le_iff_nextPotential_ge
prop:stochastic-product-tractable Full DQ.ClaimClosure.query_obstruction_boolean_corollary
prop:stochastic-sequential-bridge-fail Full DQ.ClaimClosure.posed_anchor_no_competence_no_exact_claim
prop:stochastic-sequential-strict Full DQ.StochasticSequential.stochastic_sequential_strict_separation
prop:structural-asymmetry Full DQ.ClaimClosure.SR1, DQ.ClaimClosure.SR2, DQ.ClaimClosure.SR3
@@
prop:sufficiency-char Full DQ.ClaimClosure.regime_core_claim_proved, DQ.ClaimClosure.regime_simulation_transfers_hardness
prop:temporal-equilibrium Full DQ.ClaimClosure.IE10, DQ.ClaimClosure.IE11
prop:thermo-conservation-additive Full DQ.ClaimClosure.reusable_heuristic_of_detectable
prop:thermo-hardness-bundle Full DQ.ClaimClosure.snapshot_vs_process_typed_boundary
prop:thermo-lift Full DQ.ClaimClosure.selectorSufficient_not_implies_setSufficient, DQ.ClaimClosure.separable_detectable
prop:thermo-mandatory-cost Full DQ.ClaimClosure.standard_assumption_ledger_unpack
@@
prop:typed-claim-admissibility Full DQ.ClaimClosure.sufficiency_iff_projectedOptCover_eq_opt
prop:typed-physical-transport-requirement Full DQ.Physics.ClaimTransport.physical_state_claim_of_instance_claim, DQ.Physics.ClaimTransport.physical_state_claim_of_universal_core
@@
prop:universal-solver-framing Full DQ.ClaimClosure.thermo_energy_carbon_lift_core
@@
thm:amortization Full DQ.HardnessDistribution.complete_model_dominates_after_threshold
thm:assumption-necessity Full DQ.Physics.AssumptionNecessity.physical_claim_requires_empirically_justified_physical_assumption, DQ.Physics.AssumptionNecessity.physical_claim_requires_physical_assumption
thm:bayes-from-counting Full DQ.Foundations.bayes_from_conditional, DQ.Foundations.counting_additive, DQ.Foundations.counting_nonneg, DQ.Foundations.counting_total, DQ.Foundations.entropy_contraction
thm:bayes-from-dq Full DQ.BayesOptimalityProof.bayes_is_optimal, DQ.bayesian_dq_matches_physics_dq, DQ.dq_derived_from_bayes, DQ.dq_is_bayesian_certainty_fraction
thm:bayes-optimal Full DQ.BayesOptimalityProof.KL_nonneg, DQ.BayesOptimalityProof.bayes_is_optimal, DQ.BayesOptimalityProof.crossEntropy_eq_entropy_add_KL
thm:boolean-primitive Full DQ.Physics.BoundedAcquisition.one_bit_per_transition
thm:bounded-acquisition Full DQ.Physics.BoundedAcquisition.acquisition_rate_bound
thm:bridge-boundary-represented Full DQ.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset, DQ.ClaimClosure.bounded_actions_detectable, DQ.ClaimClosure.bridge_boundary_represented_family
@@
thm:checking-duality Full DQ.checking_time_ge_witness_budget, DQ.checking_witnessing_duality_budget, DQ.no_sound_checker_below_witness_budget
thm:choice-pays Full DQ.ClaimClosure.GE3, DQ.ClaimClosure.GE7
thm:claim-integrity-meta Full DQ.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core
thm:competence-access Full DQ.ClaimClosure.IA5, DQ.ClaimClosure.IA6
thm:competence-capacity Full DQ.ClaimClosure.CH1, DQ.ClaimClosure.CH5
thm:complexity-dichotomy Full DQ.StochasticSequential.ClaimClosure.claim_tractable_subcases_to_P, DQ.StochasticSequential.complexity_dichotomy_hierarchy